Gone fishing. Debbie traveled by boat 5 miles upstream to fish in her favorite spot. Because of the 4-mph current, it took her 20 minutes longer to get there than to return. How fast will her boat go in still water?
Answer by scott8148(6628)

t = d / r ___ the upstream time is 1/3 h (20 min) longer

5 / (b - 4) = [5 / (b + 4)] + (1/3)

multiplying by 3(b-4)(b+4) ___ 15b + 60 = 15b - 60 + b^2 - 16

b^2 - 136 = 0